Calculate the PPP-adjusted GDP for each of the four countries using the information found in the table below. Instructions: Round your answers to the nearest dollar CountryGDP Ona Rye Zolfo Avon Price level PPP-adjusted GDP 10,000 12,700 14,100 23,400 -27 10 20
Answer #1

PPP-adjusted GDP formula=GDP/(1+price level)

1)

PPP-adjusted GDP (Ona)=GDP/(1+price level)=10000/(1+6%)=9434

2)

PPP-adjusted GDP (Rye)=GDP/(1+price level)=12700/(1-27%)=17397

3)

PPP-adjusted GDP (Zolfo)=GDP/(1+price level)=14100/(1-10%)=15667

4)

PPP-adjusted GDP (Avon)=GDP/(1+price level)=23400/(1+20%)=19500

the above is answer..